OpenStudy (andrejrusa123):

The solution of the equation 3d - 6 = 5 is _____.

OpenStudy (andrejrusa123):

@OswaldMurphy

OpenStudy (oswaldmurphy):

Do you want an answer or me to explain it to you?

OpenStudy (andrejrusa123):

Both if you can

OpenStudy (oswaldmurphy):

3d - 6 = 5. You want d by its self. So first add 6 to both sides. What does the equation look like then?

OpenStudy (andrejrusa123):

9-12

OpenStudy (sparkzwave12):

3D-6=5? Well you can switch D to X if it makes it simpler, but the best way to do these sorts of problems is by extracting numbers from each side to make it more even and easier to understand. 3D-6=5. Add six to both sides, making the negative six cancel out because -6+6= 0, and therefore isn't needed anymore. Now add six over to 5. That will be 11. Now the equation should look like 3D=11. Now divide each side by three. This sort of number mix up will most surely have decimals, and not be a whole number. 3D/3 is just D. 11/3 would be 3.66666667, or 3.67. In other words, D=3.67

OpenStudy (oswaldmurphy):

No, if you add 6 you will have 3d = 11. 6 + (- 6) = 0 and 6 + 5 = 11.

OpenStudy (oswaldmurphy):

Sparkz thanks but I got it. you dont want to just make a huge thing like that with the answer at then end. You try to get them to figure it out them selves. :D

OpenStudy (andrejrusa123):

the options are (3) 2/3 (1) 1/3 1/3 -1/3

OpenStudy (oswaldmurphy):

Can I finish helping you? Or do you just want the answer?

OpenStudy (andrejrusa123):

can you give me the Answer, i'm behind on a lot of stuff in school, so i'm trying get them done asap

OpenStudy (oswaldmurphy):

It would be A. 3 (2/3) = 3.6666666666

OpenStudy (yusmin):

i came up with